### v1

@@ 1 Introduction
In the context of the increasing capability of large language models (LLMs), LLM-based autonomous agents have become a new paradigm for AI applications yao2023react; xi2025rise; wang2024survey. These agents are capable of understanding instructions, invoking tools, reasoning and planning, and performing complex tasks, and are widely used in a variety of fields such as research assistants, process automation, retrieval-augmented generation, and code generation and debugging grattafiori2024llama; masterman2024landscape; qin2024tool; singh2025agentic. Although current intelligent agent systems (e.g., AutoGPT, LangChain Agents, BabyAGI)  yang2023auto; topsakal2023creating; nakajima2023task have begun to take shape, there are still numerous challenges in their generality, stability, and manageability, especially in task precision, response reliability, and system stability in situations with many tasks cemri2025multi.
Most existing agent systems rely on fixed prompts and predefined tool-use workflows, limiting their ability to understand task intent, select tools dynamically, and manage context effectively e2025rag; qin2024tool; schick2023toolformer. Numerous works have shown that dynamic prompt, tool retrieval, and memory management play a key role in optimizing agent execution. For instance, MCP-Zero fei2025mcp enables active tool discovery for unseen tasks, while multi-agent designs zhou2025multi enhance planning and coordination through optimized prompts and collaboration strategies. The Model Context Protocol (MCP) hou2025model standardizes context exchange, ensuring reliable tool invocation across systems. Together, these advances underscore that robust prompting, intelligent tool access, and effective memory are essential for high-quality agent execution.
@@
Rather than evaluating isolated modules, this paper builds a unified framework named Jenius-Agent, which improves task accuracy, efficiency, and contextual robustness, aligned with emerging agent communication protocols (e.g., MCP, ACP, A2A) ibm2024acp; google2024a2a.
@@ 2 Related Work
With the rapid advancement of LLMs, research on agent systems has expanded in multiple directions. A typical ReAct yao2022react style autonomous agent workflow, illustrated in Figure 1, follows an interleaved cycle of reasoning, action, and feedback. In this paradigm, the LLM is no longer a passive text generator but serves as a central planner. This iterative planner–tool–environment interaction highlights three optimization focuses:(i) prompt optimization for reliable reasoning, (ii) tool selection for precise external interactions, and (iii) memory management to preserve context and limit token usage.
@@ 2.1 Prompt Engineering and Optimization
DSPy khattab2024dspy provides a modular framework for prompt compilation, improving clarity and debuggability, while Prompt Engineering in LLMs marvin2023prompt summarizes practical design principles. Reflect-Retry-Reward bensal2025reflect further introduces adaptive optimization through self-reflection and reinforcement learning. However, current approaches remain unreliable in tool-based or long-horizon tasks, as they often depend on static templates, lack systematic evaluation, and struggle to control structured output or tool invocation consistency.
@@ 2.2 Tool Selection and Invocation
Recent efforts such as MCP-Zero fei2025mcp explore adaptive tool discovery using semantic embeddings, and BioMedTools liu2025biomedtools illustrates domain-specific coordination via curated tool metadata.However, prior approaches frequently assume clean tool descriptions and fixed invocation formats, whereas real-world systems face noisy metadata, semantic mismatches, and poorly handled runtime errors, resulting in unstable tool retrieval and parameter usage.
@@ 2.3 Memory Management
StateFlow wu2024stateflow and ReAct yao2023react demonstrate state-aware and reasoning-aligned memory design, while Recursively Summarizing wang2025recursively introduces hierarchical compression for long-term coherence.Existing methods often use simple window-based truncation or stage-level summaries, risking loss of critical early context or fine-grained tool dependencies, lacking tight integration between reasoning and memory retention.
@@ 3.2 Context-Aware Tool Orchestration
All tools are represented as high-dimensional embeddings using the Qwen3 Embedding model zhang2025qwen3, and tool selection is formalized as a semantic similarity ranking process. The similarity between the query and each tool embedding determines their conceptual proximity, which serves as the basis for ranking and subsequent relevance filtering.
@@ 4.2.1 APIGen: A Single-Turn Tool-Use Benchmark
APIGen liu2024apigen evaluates models’ ability to interpret natural language queries into structured API calls. It contains 60K annotated samples across 21 categories (3.6K APIs, 2.3 parameters per call). Each instance includes a user query, candidate APIs, and a ground-truth function call.

### v2

@@ Abstract
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
In the context of the increasing capability of large language models (LLMs), LLM-based autonomous agents have become a new paradigm for AI applications Yao et al. (2023); Xi et al. (2025); Wang et al. (2024). These agents are capable of understanding instructions, invoking tools, reasoning and planning, and performing complex tasks, and are widely used in a variety of fields such as research assistants, process automation, retrieval-augmented generation, and code generation and debugging Dubey et al. (2024); Masterman et al. (2024); Qin et al. (2024); Singh et al. (2025). Although current intelligent agent systems (e.g., AutoGPT, LangChain Agents, BabyAGI)  Yang et al. (2023); Topsakal and Akinci (2023); Nakajima (2023) have begun to take shape, there are still numerous challenges in their generality, stability, and manageability, especially in task precision, response reliability, and system stability in situations with many tasks Cemri et al. (2025).
Most existing agent systems rely on fixed prompts and predefined tool-use workflows, limiting their ability to understand task intent, select tools dynamically, and manage context effectively e Aquino et al. (2025); Qin et al. (2024); Schick et al. (2023). Numerous works have shown that dynamic prompt, tool retrieval, and memory management play a key role in optimizing agent execution. For instance, MCP-Zero Fei et al. (2025) enables active tool discovery for unseen tasks, while multi-agent designs Zhou et al. (2025) enhance planning and coordination through optimized prompts and collaboration strategies. The Model Context Protocol (MCP) Hou et al. (2025) standardizes context exchange, ensuring reliable tool invocation across systems. Together, these advances underscore that robust prompting, intelligent tool access, and effective memory are essential for high-quality agent execution.
@@
Rather than evaluating isolated modules, this paper builds a unified framework named Jenius-Agent, which improves task accuracy, efficiency, and contextual robustness, aligned with emerging agent communication protocols (e.g., MCP, ACP, A2A) IBM BeeAI (2024); Google (2024).
@@ 2 Related Work
With the rapid advancement of LLMs, research on agent systems has expanded in multiple directions. A typical ReAct Yao et al. (2022) style autonomous agent workflow, illustrated in Figure 1, follows an interleaved cycle of reasoning, action, and feedback. In this paradigm, the LLM is no longer a passive text generator but serves as a central planner. This iterative planner–tool–environment interaction highlights three optimization focuses:(i) prompt optimization for reliable reasoning, (ii) tool selection for precise external interactions, and (iii) memory management to preserve context and limit token usage.
@@ 2.1 Prompt Engineering and Optimization
DSPy Khattab et al. provides a modular framework for prompt compilation, improving clarity and debuggability, while Prompt Engineering in LLMs Marvin et al. (2023) summarizes practical design principles. Reflect-Retry-Reward Bensal et al. (2025) further introduces adaptive optimization through self-reflection and reinforcement learning. However, current approaches remain unreliable in tool-based or long-horizon tasks, as they often depend on static templates, lack systematic evaluation, and struggle to control structured output or tool invocation consistency.
@@ 2.2 Tool Selection and Invocation
Recent efforts such as MCP-Zero Fei et al. (2025) explore adaptive tool discovery using semantic embeddings, and BioMedTools Liu et al. (2025) illustrates domain-specific coordination via curated tool metadata.However, prior approaches frequently assume clean tool descriptions and fixed invocation formats, whereas real-world systems face noisy metadata, semantic mismatches, and poorly handled runtime errors, resulting in unstable tool retrieval and parameter usage.
@@ 2.3 Memory Management
StateFlow Wu et al. and ReAct Yao et al. (2023) demonstrate state-aware and reasoning-aligned memory design, while Recursively Summarizing Wang et al. (2025) introduces hierarchical compression for long-term coherence.Existing methods often use simple window-based truncation or stage-level summaries, risking loss of critical early context or fine-grained tool dependencies, lacking tight integration between reasoning and memory retention.
@@ 3.2 Context-Aware Tool Orchestration
All tools are represented as high-dimensional embeddings using the Qwen3 Embedding model Zhang et al. (2025), and tool selection is formalized as a semantic similarity ranking process. The similarity between the query and each tool embedding determines their conceptual proximity, which serves as the basis for ranking and subsequent relevance filtering.
@@ 4.2.1 APIGen: A Single-Turn Tool-Use Benchmark
APIGen Liu et al. (2024) evaluates models’ ability to interpret natural language queries into structured API calls. It contains 60K annotated samples across 21 categories (3.6K APIs, 2.3 parameters per call). Each instance includes a user query, candidate APIs, and a ground-truth function call.